Circular No. 6773

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dr IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (R = V = 18.7) on V and R CCD images taken on Nov. 7 by S. Chan at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997dr is located near Abell 3112 at α = 3^{h}18^{m}49^{s}.21, δ = −44^{o}28′13′′.4 (equinox 2000.0), only 0′′.5 north and 0′′.2 east of its host galaxy. A nearby star has position end figures 55^{s}.51, 28′′.3. SN 1997dr was not visible on similar images taken on Oct. 17 (limiting mag R = 21), but is present in V and R images taken at the Mount Stromlo 0.76-m telescope on Nov. 17, when it appears 0.2 mag brighter. Finding charts are available, as usual, at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/. N. Suntzeff re-

ports that images obtained on Nov. 17 UT with the Cerro Tololo 0.9-m telescope show that photometric identification is difficult due to the supernova’s proximity to the galaxy nucleus.

SUPERNOVA 1997ds IN MCG –01-57-007 Y. L. Qiu, Q. Y. Qiao, W. D. Li, and J. Y. Hu, Bejing Astronomical

Observatory (BAO); and A. Esamdin, Urumqi Astronomical Station, report their discovery of an apparent supernova (mag ∼ 16.3) on unfiltered CCD images taken with a 0.6-m telescope on Nov. 19.49 UT. SN 1997ds is located at α = 22^{h}24^{m}11^{s}.51, δ = −3^{o}29′10′′.5 (equinox 2000.0), which is 0′′.9 west and 14′′.2 south of the center of MCG –01-57-007. The supernova was confirmed on images taken on Nov. 17.47 (mag ∼ 16.2). Unfiltered CCD images of the same field taken on Nov. 4 show no star at this location.

EF PEGASI Visual magnitude estimates by P. Schmeer, Bischmisheim, Germany

(cf. IAUC 6767): Nov. 3.735 UT, 10.9; 3.883, 11.0; 7.753, 11.4; 8.728, 11.8; 9.765, 11.9; 13.734, 12.4.

Corrigendum. On IAUC 6767, last line, for p.a. 305^{o}. read p.a. 112^{o}.

COMET C/1997 T1 (UTSUNOMIYA) Visual m_{1} estimates: Nov. 1.03 UT, 10.3 (G. W. Kronk, Troy, IL,

0.33-m reflector); 6.85, 9.9 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 18.98, 9.9 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 19.74, 10.4 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector).

1997 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green